OpenStudy (anonymous):

Plot Three Points For The Equation: 5x-4y=-18

OpenStudy (anonymous):

Example: 5(2)-4y=18 10-4y=18 -4y=-28 y=28 (x=2,y=7) Answer: (2,7) (-2,2) (-6,-3)